OpenStudy (anonymous):

Express the complex number in trigonometric form. -4

OpenStudy (anonymous):

\[-4=r \left( \cos \theta+\iota \sin \theta \right)\] equating real and imaginary parts \[r \cos \theta=-4,r \sin \theta=0\] square and divide \[r^2\left( \cos ^2\theta+\sin ^2\theta \right)=(-4)^2+0^2=16\] \[r^2=16,r=4\] \[r \cos \theta=-4,4 \cos \theta=-4,\cos \theta=-1=\cos \pi ,\theta=\pi \] \[Hence -4=4 \cos \pi+ \iota 4 \sin \pi \]